OpenStudy (anonymous):

Use differentials (or equivalently, a linear approximation ) to approximate ln(1.01) as follows: Let f(x)=ln(x) and find the equation of the tangent line to f(x) at a "nice" point near 1.01. Then use this to approximate ln(1.01). Approximation=

OpenStudy (anonymous):

\[f(x) = \ln(x)\] \[\frac{df}{dx} = \frac{1}{x} \] So, the equation of the tangent line at x = 1 has slope \[\frac{df}{dx} = \frac{1}{1} = 1\] The equation of that line is \[(y-0) = 1(x-1)\rightarrow y = x - 1 \] So, near x = 1, we have that \[f(x) \approx x - 1\] so, \[\ln(1.01) \approx (1.01) - 1 = 0.01\]

OpenStudy (anonymous):

Thank you so much!